Robert is 2 years older than Jazmine. Jazmine is 3 years older than Chris, who is one year less than 1/2 of Michael's age. Together, the sum of their ages is 30. How old is each child? Show your work and explain your reasoning
Answer by lwsshak3(11628)

Robert is 2 years older than Jazmine. Jazmine is 3 years older than Chris, who is one year less than 1/2 of Michael's age. Together, the sum of their ages is 30. How old is each child? Show your work and explain your reasoning
***
Michael's age=x
Chris' age=(x/2)-1
Jasmine's age=(x/2)-1+3=(x/2)+2
Robert's age=(x/2)+2+2=(x/2)+4
sum of their ages=30
x+(x/2)-1+(x/2)+2+(x/2)+4=30
(5x/2)+5=30
5x/2=25
5x=50
x=10
(x/2)-1=4
(x/2)+2=7
(x/2)+4=9
Michael's age=10
Chris' age=4
Jasmine's age=7
Robert's age=9